OpenStudy (anonymous):

Let f(x) = (1-2x+x^2)/(abs(1-x)) find lim x--> 1+ and lim x-->1-

myininaya (myininaya):

factor top remeber |1-x|=1-x if 1-x>0=>1>x |1-x|=-(1-x)= if 1-x<0=> 1<x

OpenStudy (anonymous):

after all that please tell me i didn't mess it up!

OpenStudy (anonymous):

oh yes i did !!!!!

OpenStudy (anonymous):

after all that ... i must be tired. numerator is \[(1-x)^2\] not \[(1-x)(1+x)\]

OpenStudy (anonymous):

is the answer 1 for both? thats what i got

OpenStudy (zarkon):

no

OpenStudy (anonymous):

can you help zarkon?

OpenStudy (zarkon):

\[\frac{(1-x)^2}{|1-x|}\] \[\frac{1-x}{|1-x|}(1-x)\] taking limit we get -1*0=0 or 1*0=0

OpenStudy (anonymous):

it is still true that \[\frac{1-x}{ |1-x|} = \left\{\begin{array}{rcc} 1 & \text{if} & x <1 \\ -1& \text{if} & x >1 \end{array} \right. \] so you have \[f(x) =\frac{(1-x)(1-x)}{|1-x|}= \left\{\begin{array}{rcc} 1-x & \text{if} & x <1 \\ x-1& \text{if} & x>1 \end{array} \right. \]

OpenStudy (anonymous):

so (1-x) cancle with denominator and left with numerator 1-x right

OpenStudy (anonymous):

is it continuous at x = 1?

OpenStudy (zarkon):

no..it is not defined at 1

OpenStudy (anonymous):

thanks